8. The cesium iodide (CsI) molecule has an atomic separation of 0.127 nm. (a) Determine the energy of the second excited rotational state, with J 5 2. (b) Find the frequency of the photon absorbed in the J 5 1 to J 5 2 transition.
Answer #1


I comes Centre of Moss k = 0.1271 Moment of Wertats about yans I = (132-9 * 114.8 x 1066*10 + (0-127) x 10 132.9 + 114.8. I

Ephoton = 4 Enot E, -E;-) = $? [JCI+) - (1-) Ephoton hf Hot Not = Putting J = 2 YZ2 I = 2h u2I 2* 6-626*104 473/1.63*10-45) =